Use the given information to find the number of degrees of​ freedom, the critical values chi Subscript Upper L Superscript 2 and chi Subscript Upper R Superscript 2​, and the confidence interval estimate of sigma. It is reasonable to assume that a simple random sample has been selected from a population with a normal distribution. Nicotine in menthol cigarettes 80​% ​confidence; n=25​, s=0.28 mg.
Answer #1

Solution :

Given that,

c = 0.80

s = 0.28

n = 25

At 80% confidence level the \chi ^{2} is ,

\alpha = 1 - 80% = 1 - 0.80 = 0.20

\alpha / 2 = 0.20 / 2 = 0.10

\chi ^{2}\alpha /2,df = \chi ^{2} 0.10,24 = 33.20

and

\chi ^{2}1- \alpha /2,df = \chi ^{2} 0.90,24 = 15.66

\chi2L = \chi 2\alpha/2,df = 33.20

\chi2R = \chi 21 - \alpha /2,df = 15.66

The 80% confidence interval for \sigma is,

s \sqrt{} (n-1) / \chi ^{2} \alpha /2,df < \sigma < s \sqrt{} (n-1) / \chi ^{2} 1- \alpha /2,df

0.28 \sqrt{} ( 25 - 1 ) / 33.20 < \sigma < 0.28 \sqrt{} ( 25 - 1 ) / 15.66

0.24 < \sigma < 0.35

( 0.24 , 0.35)